OpenStudy (anonymous):

Mary Kim invested $964 in 18 shares of stock. A month later, she sold the 18 shares at $61.48 per share. She also invested in 350 shares of stock for a total of $27, 422.50. She sold this stock for $85.35 a share and paid $495 in total commissions. What was Mary Kim’s profit or loss on the transactions to the nearest dollar?

OpenStudy (callisto):

18*61.45+85.35*(350) - 964-27422.5-495 = 30978.6-27891.5 = $3087

OpenStudy (callisto):

profit

OpenStudy (anonymous):

Let us take the first investment of $964 When she sold her shares for $61.48 each, she earned $1106.64 Net profit on this transaction is $142.64 Let us now look at the second investment of $27, 422.50 Money she made by selling at $ 85.35 = 350* 85.35 = 29872.5 So profit is$29, 872.5 - $27, 422.5 - $495 = $1955 Therefore she profited by $2097.64

sam (.sam.):

18*61.48 - 964 = 142.64 total profit 350*85.35 - 27422.50= 2450 total profit Total profit: 2450.00 + 142.64 - 495.00 = $2097.64